OpenStudy (emilyjones284):

MEDAL AND FAN! (: What is the explicit formula for the sequence? 3, 1, –1, –3, –5, . . .

OpenStudy (anonymous):

x-2

OpenStudy (anonymous):

3-2=1 , 1-2=-1, -1-2=-3

OpenStudy (emilyjones284):

Thats not an answer choice

OpenStudy (emilyjones284):